Three copper wires have their lengths in the ratio 5:3:1 and their masses are in the ratio 1:3:5. Their electrical resistance will be in the ratio

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
l1:l2:l3=5:3:1 
m1:m2:m3=1:3:5 
thickness will be same of each of the 3 since material is same. 
m1/v1 =m2/v2 =m3/v3 
=>1/a1l1 = 3/a2l2 = 5/a3l3 
=>a1:a2:a3=1:5:25 
R1:R2:R3= l1/a1 : l2/a2 : l3/a3 
=5x/y : 3x/5y : x/25y 
=5:3/5:1/25 
=125:15:1 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)